OpenStudy (anonymous):

Just another very easy problem: In an exam, \(5%\) applicants were ineligible and \(85%\) of eligible belonged to A category. If \(4275\) eligible belong to other categories, how many applied? If you google you will be spoiling your fun :)

OpenStudy (anonymous):

30000

OpenStudy (anonymous):

Apparently that is not the correct answer.

OpenStudy (anonymous):

dont laugh but i think its 4,365

OpenStudy (anonymous):

Nopes :)

OpenStudy (anonymous):

i got it now we'll suppose no of total applicants = x 95% of which are eligible 5% are ineligible so,

OpenStudy (anonymous):

95/100 * x is the eligible category

OpenStudy (anonymous):

85/100 * x are category A and eligible 100-85 = 15 so 15/100*x are other category and eligible so, 15% of (95% of x) so 15/100 * 95/100* x 0.1425x but the question tells us that other category guyz are 4275 so 0.1425x = 4275 x= 4275/0.1425 x = 30,000 @FoolForMath i solved it :)

OpenStudy (anonymous):

hey i won :)

OpenStudy (anonymous):

I got the same answer. But to some weird reason the answer is 35,000. You may Google now :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

any particular site ?

OpenStudy (anonymous):

@FoolForMath then i was correct u told i was wrong hm.......!!

OpenStudy (anonymous):

wow thanks api

OpenStudy (lalaly):

hehe :D:D

OpenStudy (lalaly):

@FoolForMath